A cleaned, dried aluminum can (Cherry Pepsi) was recently weighed. The recorded mass was 13.2518 grams. Assuming that 0.5000 g of the weight was other materials (paint, aluminum oxide), what would the theoretical yield be (on a mass basis) if all of the aluminum in the can was converted into alum? Remember to report your result with the correct units and precision.
Answer #1

Overall reactuin

K+(aq) + Al3+(aq) +2SO42-(aq) +12H2O(l) ----->KAl(SO4)2·12H2O(s)

1 mole of Al gives 1 mole of KAl(SO4)2. 12H2O

mass of Al in the sample = 13.2518-0.5000=12.7518 gm

moles of Al = mass/ molar mass = 12.7518/27=0.4723

moles of Alum formed , KAl (SO4)2. 12H2O = 0.4723,

molar mass of Alum = 474, mass of alum = 474*0.4723 gm =223.8 gm ( theoretical yield)